the voltage across diode at room temperature of 300 k is 0.71v when 2.5 ma current flows through it .if the voltage increases to 0.8v .calculate the new diode current
1
Expert's answer
2021-08-24T14:49:08-0400

∆v=(0.8-0.71)=0.09v

T=300k

"I_0=2.5mA"

"I=I_0e^\frac{e∆V}{KT}"

Put value


"I=2.5\times10^{-3}(e^\frac{1.6\times10^{-19}\times0.09}{2\times1.38\times10^{-23}\times300}-1)"

"I=2.5\times10^{-3}(e^{1.76}-1)\\I=2.5\times10^{-3}(5.69-1)=11.7mA"